# Flipped rotating axion non-minimally coupled to gravity: Baryogenesis and Dark Matter

### v1

@@ Abstract
We demonstrate that the co-genesis of baryon asymmetry and dark matter can be achieved through the rotation of an axion-like particle, driven by a flip in the vacuum manifold’s direction at the end of inflation. This can occur if the axion has a periodic non-minimal coupling to gravity, while preserving the discrete shift symmetry. In non-oscillating inflation models, after inflation there is typically a period of kination (with $w=1$). In this case, it is shown that the vacuum manifold of the axion is flipped and the axion begins rotating in field space, because it can slide across the decreasing potential barrier as in Ricci reheating. Such a rotating axion can generate the baryon asymmetry of the Universe through spontaneous baryogenesis, while at later epochs it can oscillate as dark matter. The period of kination makes the primordial gravitational waves (GW) generated during inflation sharply blue-tilted which constrains the parameter space due to GW overproduction, while being testable by next generation CMB experiments. As a concrete example, we show that such a cogenesis of baryon asymmetry and dark matter can be realized for the axion as the Majoron in the Type-I seesaw setup, predicting mass ranges for the Majoron below sub eVs, with right-handed neutrino mass above $\mathcal{O}(10^{8})$ GeV. We also show that in order to avoid fragmentation of the axion condensate during the rotation, we require the non-minimal coupling $\xi\sim(f/m_{P})^{2}$ or somewhat larger, where $f$ is the axion decay constant.
@@ I Introduction
Cosmic inflation is a period of exponential increase in the size of the Universe which accounts for the otherwise fine-tuned initial conditions of the hot Big Bang history, and is responsible for generating the primordial density perturbations seeding structure formation Starobinsky (1980); Sato (1981); Kazanas (1980); Guth (1981). The measurements of the cosmic microwave background (CMB) radiation, help us unravel the microscopic model of inflation which come in many avatars. The recent CMB observations from the Planck satellite mission Aghanim et al. (2020) have led to severe constraints on several inflationary models but neither it has been able to yet pin down upon a specific scenario nor it has gives us definitive insights into embedding of the inflationary paradigm into fundamental particle physics theory like that of Grand Unification (GUT). Upcoming CMB missions like the Simons Observatory Ade et al. (2019), the LiteBIRD satellite Hazumi et al. (2019); Sugai et al. (2020) or the ground-based CMB Stage 4 program Abazajian et al. (2022), Ali CMB Polarization Telescope Li et al. (2019) and CMB-Bharat  Adak et al. (2022) will be able to further investigate on this direction, particularly so concerning the measurement of the tensor-to-scalar ratio $r$ Martin et al. (2014), which basically depicts the energy scale of inflation which is determined once the BB-mode correlations are measured and dark radiation measurements $\Delta N_{\rm eff}$. This leads to important science targets to go beyond this paradigm and also test the indirect evidence of such inflationary characteristics like for instance, observables signatures like primordial non-gaussianity which may point towards inflaton interactions or the microscopic details of the reheating epoch in the post-inflationary epoch Bassett et al. (2006); Martin and Ringeval (2010); Martin et al. (2015); Meerburg et al. (2019); Allahverdi et al. (2010); Amin et al. (2014).
@@
In scenarios with the inflationary paradigm having non-oscillatory quintessential potential for the inflaton, rather generically, in the post-inflationary era the universe becomes dominated by the kinetic-energy of the inflaton. This period is called kination Joyce and Prokopec (1998); Gouttenoire et al. (2021a). The essential idea of Ricci reheating in this framework is that, any spectator scalar field non-minimally coupled to gravity undergoes a second-order phase transition during the time when the Universe transitions from an inflating background to a kination dominated background when the Ricci scalar changes sign and turns negative. In Ricci reheating, after the phase transition, the non-minimally coupled scalar field undergoes coherent oscillations, which amount to particles that decay and reheat the Universe Dimopoulos and Markkanen (2018); Opferkuch et al. (2019); Bettoni et al. (2022); Laverda and Rubio (2024). In this paper, the above phase transition triggers a rotation in field space.
@@
In the most generic context there exists several motivations of having spontaneously broken global symmetries involving solutions to a myriad of problems in the Standard Model (SM) of particle physics, to name a few, the Peccei-Quinn (PQ) symmetry Peccei and Quinn (1977) as a solution to the strong CP problem, lepton number or baryon number symmetry Chikashige et al. (1981) in the context with the microscopic origin the SM neutrino masses, and flavor symmetries to explain the structure of CKM quark or PNMS neutrino mass matrices as explanation of the observed pattern of the fermion masses and mixings Froggatt and Nielsen (1979). Nonetheless breaking of any such symmetries will inevitably lead the existence of very light degrees of freedom, known as Nambu-Goldstone bosons (NGB) which have different names in different physics contexts, namely the QCD axion Weinberg (1978); Wilczek (1978) for the PQ symmetry, a Majoron for lepton number Chikashige et al. (1981), and a familon or flavon Reiss (1982); Wilczek (1982) for the case involving flavor symmetry. For the purpose of this paper, we simply focus on any such pseudo Nambu-Goldstone boson or axion-like particle (ALP) generically but we dub it as “axion” for the sake of simplicity. The axions we study can even have a fundamental origin in the context of the string axiverse and can be present in more than avatars Arvanitaki et al. (2010).
@@
The common lore is that the axion field is initially static and misaligned at a non-zero field value, and later oscillates when its mass becomes comparable to the Hubble expansion rate of the Universe. However, the axion field may instead be initially rotating in the complex field space Gouttenoire et al. (2021a); Li et al. (2014); Co et al. (2022a); Gouttenoire et al. (2021b); Harigaya et al. (2023a, b); Chung and Tadepalli (2024); Duval et al. (2024). Such an initial kick in the angular direction is usually realized at high-radial field values, through operators explicitly breaking the global symmetry. These dynamics at high radial field values have been crucial in cosmology, for e.g., in realizing Affleck-Dine baryogenesis Affleck and Dine (1985). Such setups have been explored further in the context of baryogenesis Co and Harigaya (2020), for modifying the dark matter abundance through the kinetic misalignment Co et al. (2020a, b), for cogenesis of both dark matter and the baryon asymmetry Co et al. (2021a, b) and also in several other related contexts Co et al. (2021c); Kawamura and Raby (2022); Co et al. (2021d, 2022b); Barnes et al. (2023); Co et al. (2023); Badziak and Harigaya (2023); Berbig (2024); Chun and Jung (2024); Chao and Peng (2023); Barnes et al. (2024); Datta et al. (2024).
@@
In this paper, we explore the possibility of generating such a rotation of the axion, without delving into the above route, such that the radial mode still remains at the minima of its potential. This can take place if the height of the axion potential barrier keeps on decreasing, such that the axion can slide over,in a similar manner to Ricci reheating Opferkuch et al. (2019); Bettoni et al. (2022) (see also Refs. Figueroa and Byrnes (2017); Nakama and Yokoyama (2019)). Such an alternative was explored recently in Chun et al. (2024), with a time-dependent axion decay constant, realized through symmetry non-restoration. Here, we propose a scenario where an axion couples non-minimally to gravity, in a periodic form, thereby respecting the shift-symmetry of any pNGB action Ferreira et al. (2018). In contrast to Refs. Gouttenoire et al. (2021a); Li et al. (2014); Co et al. (2022a); Gouttenoire et al. (2021b); Harigaya et al. (2023a); Duval et al. (2024), our rotating axion always remains a spectator field. Because of its non-minimal coupling to gravity which depends on the cosmological background through the Ricci scalar $R$, the axion vacuum manifold evolves depending on the barotropic parameter $w$. While $w$ changes from $-1$ during inflation to $+1$ during kination (right after inflation), the axion vacuum manifold flips in the opposite direction, generating an initial kinetic energy density for the axion. Interestingly, if the Universe happens to be kination-dominated after inflation, the potential barrier height being determined by the cosmological background, redshifts and decreases fast enough such that the axion can keep sliding across the barrier, and hence rotate in the complex field space.
@@
We utilise such a dynamics of the axion field to successfully generate both the baryon asymmetry and the dark matter abundance, dubbed as cogenesis. While the baryon asymmetry can be generated during the rotation through spontaneous baryogenesis Cohen and Kaplan (1987, 1988), the dark matter abundance can arise at a much later epoch, after the bare mass potential of the axion starts to dominate over its kinetic energy density. The observed value of baryon asymmetry constrains the non-minimal coupling $\xi$, while the dark matter abundance relates the axion decay constant to its mass. Importantly, because of the kination-dominated stiff era, the spectrum of the primordial gravitational waves (GWs) generated during inflation becomes blue-tilted for modes entering the horizon during kination. Such an increase in the GW amplitude tightly constrains our parameter space, because of excess GW during Big Bang Nucleosynthesis (BBN).
@@
Throughout the paper we use natural units for which $\hbar=k_{B}=c=1$ and $8\pi G=m_{P}^{-2}$, with $m_{P}=2.43\times 10^{18}\,$eV being the reduced Planck mass. The signature of the metric is positive.
@@ II The Framework
first introduced in Ref. Ferreira et al. (2018). In the above, $(\partial\phi)^{2}\equiv\partial_{\mu}\phi\,\partial^{\mu}\phi$ and
@@
When $|\phi|\ll f$, $[1-\cos(\phi/f)]\simeq\frac{1}{2}(\phi/f)^{2}$. Then, the theory in Eq. (1) becomes
@@
which is reminiscent of the usual Lagrangian density of a scalar field with a non-minimal coupling to gravity. However, for a rotating axion the approximation $|\phi|\ll f$ is not valid ($-\pi f<\phi\leq\pi f$), and we have to consider the full Lagrangian density in Eq. (1) without this approximation.
@@
In FRW spacetime $R=3(1-3w)H^{2}$, where $w$ and $H$ are the barotropic and the Hubble parameter, respectively. We consider that our axion is a spectator field in a non-oscillatory inflationary scenario, where inflation (with $w=-1$) is followed by a period of kination (with $w=1$). We see, therefore, that during inflation, the axion effective potential is
@@
For the rolling of the axion to occur we need that $||V_{\rm eff}^{\prime\prime}||>(\frac{3}{2}H)^{2}$, where $||V_{\rm eff}^{\prime\prime}||=3\xi m_{P}^{2}H^{2}/f^{2}$. This condition suggests $\xi>\frac{3}{4}(f/m_{P})^{2}$, such that the range of $\xi$ is
@@
This condition also makes sure that during inflation the axion is driven towards the minimum expectation value $\phi=\pi\,f$. Because the axion is heavy during inflation, it does not undergo particle production. As a result, it does not introduce isocurvature perturbation.
@@
As we have assumed that $M$ is very small, we expect that the effective axion mass is $m_{\phi}^{2}=||V^{\prime\prime}(\phi)||\simeq M^{4}/f^{2}\ll H^{2}$ after reheating. This means that the rotating axion eventually freezes with $\phi={\cal O}(f)$, until the decreasing $H(t)=1/2t$ catches up with $m_{\phi}$. At that point, the axion unfreezes and begins coherent oscillations in its potential, which soon becomes quadratic with mass $m_{\phi}=M^{2}/f$. The energy density of the oscillating axion condensate decreases as pressureless matter such that $\rho_{\phi}\propto a^{-3}$ from now on. As a result, the axion condensate eventually dominates the radiation background.
@@
where $0\lesssim\lambda\lesssim 1$ is its self-coupling constant. To the above we add a temperature correction due to interaction with the thermal bath
@@
Assuming that the thermal bath is generated at the end of inflation (there are many mechanisms that do so naturally, e.g., by instant preheating Felder et al. (1999b); Dimopoulos et al. (2018) or Ricci reheating Dimopoulos and Markkanen (2018); Opferkuch et al. (2019); Bettoni et al. (2022)), we obtain an estimate of $T_{\rm max}$ in the following way. After the end of inflation there is a period of kination, when the density of the Universe scales as $\rho\propto a^{-6}$, while radiation, which appears at the end of inflation, scales as $\rho_{r}\propto a^{-4}$. This means that
@@
This bound ensures that $|\Phi|\neq 0$ and the axion always exists. Considering GUT-scale inflation $H_{\rm end}\sim 10^{-5}m_{P}$. Then, because $T_{\rm reh}\gtrsim 10^{7}\,$eV to avoid overproduction of gravitational radiation at BBN (cf. Eq. (47)), we obtain

### v2

@@ Abstract
We demonstrate that the co-genesis of baryon asymmetry and dark matter can be achieved through the rotation of an axion-like particle (but not the QCD axion), driven by a flip in the vacuum manifold’s direction at the end of inflation. This can occur if the axion has a periodic non-minimal coupling to gravity, while preserving the discrete shift symmetry. In non-oscillating inflation models, after inflation there is typically a period of kination (with $w=1$). In this case, it is shown that the vacuum manifold of the axion is flipped and the axion begins rotating in field space, because it can slide across the decreasing potential barrier as in Ricci reheating. Such a rotating axion can generate the baryon asymmetry of the Universe through spontaneous baryogenesis, while at later epochs it can oscillate as dark matter. The period of kination makes the primordial gravitational waves (GW) generated during inflation sharply blue-tilted which constrains the parameter space due to GW overproduction, while being testable by next generation CMB experiments. As a concrete example, we show that such a cogenesis of baryon asymmetry and dark matter can be realized for the axion as the Majoron in the Type-I seesaw setup, predicting mass ranges for the Majoron below sub eVs, with right-handed neutrino mass above $\mathcal{O}(10^{8})$ GeV. We also show that in order to avoid fragmentation of the axion condensate during the rotation, we require the non-minimal coupling $\xi\sim(f/m_{P})^{2}$ or somewhat larger, where $f$ is the axion decay constant.
@@ I Introduction
Cosmic inflation is a period of exponential increase in the size of the Universe which accounts for the otherwise fine-tuned initial conditions of the hot Big Bang history, and is responsible for generating the primordial density perturbations seeding structure formation Starobinsky (1980); Sato (1981); Kazanas (1980); Guth (1981). The measurements of the cosmic microwave background (CMB) radiation, help us unravel the microscopic model of inflation which come in many avatars. The recent CMB observations from the Planck satellite mission Aghanim et al. (2020) have led to severe constraints on several inflationary models but neither it has been able to yet pin down upon a specific scenario nor it has given us definitive insights into embedding of the inflationary paradigm into fundamental particle physics theory like that of Grand Unification (GUT). Upcoming CMB missions like the Simons Observatory Ade et al. (2019), the LiteBIRD satellite Hazumi et al. (2019); Sugai et al. (2020) or the ground-based CMB Stage 4 program Abazajian et al. (2022), Ali CMB Polarization Telescope Li et al. (2019) and CMB-Bharat  Adak et al. (2022) will be able to further investigate on this direction, particularly so concerning the measurement of the tensor-to-scalar ratio $r$ Martin et al. (2014), which basically depicts the energy scale of inflation which is determined once the BB-mode correlations are measured and dark radiation measurements of $\Delta N_{\rm eff}$. This leads to important science targets to go beyond this paradigm and also test the indirect evidence of such inflationary characteristics like for instance, observable signatures like primordial non-gaussianity which may point towards inflaton interactions or the microscopic details of the reheating epoch in the post-inflationary epoch Bassett et al. (2006); Martin and Ringeval (2010); Martin et al. (2015); Meerburg et al. (2019); Allahverdi et al. (2010); Amin et al. (2014).
@@
In scenarios with the inflationary paradigm having non-oscillatory quintessential potential for the inflaton, rather generically in the post-inflationary era, the Universe becomes dominated by the kinetic energy density of the inflaton. This period is called kination Joyce and Prokopec (1998); Gouttenoire et al. (2021a). The essential idea of Ricci reheating in this framework is that, any spectator scalar field non-minimally coupled to gravity undergoes a second-order phase transition during the time when the Universe transitions from an inflating background to a kination dominated background when the Ricci scalar changes sign and turns negative. In Ricci reheating, after the phase transition, the non-minimally coupled scalar field undergoes coherent oscillations, which amount to particles that decay and reheat the Universe Dimopoulos and Markkanen (2018); Opferkuch et al. (2019); Bettoni et al. (2022); Laverda and Rubio (2024). In this paper, the above phase transition triggers a rotation in field space.
@@
In the most generic context there exists several motivations of having spontaneously broken global symmetries involving solutions to a myriad of problems in the Standard Model (SM) of particle physics, to name a few, the Peccei-Quinn (PQ) symmetry Peccei and Quinn (1977) as a solution to the strong CP problem, lepton number or baryon number symmetry Chikashige et al. (1981) in the context with the microscopic origin the SM neutrino masses, and flavor symmetries to explain the structure of CKM quark or PNMS neutrino mass matrices as explanation of the observed pattern of the fermion masses and mixings Froggatt and Nielsen (1979). Nonetheless, the breaking of any such symmetries inevitably leads to the existence of very light degrees of freedom, known as Nambu-Goldstone bosons (NGB) which have different names in different physics contexts, namely the QCD axion Weinberg (1978); Wilczek (1978) for the PQ symmetry, a Majoron for lepton number Chikashige et al. (1981), and a familon or flavon Davidson and Wali (1982); Reiss (1982); Wilczek (1982); Davidson et al. (1984a, b) for the case involving flavor symmetry. For the purpose of this paper, we simply focus on any such pseudo NGB or axion-like particle (ALP) generically but we dub it as “axion” for the sake of simplicity (to avoid thermal corrections, we do not consider the QCD axion). The axions we study can even have a fundamental origin in the context of the string axiverse and can be present in more avatars Arvanitaki et al. (2010).
Such axion-like particles can couple to the Standard Model through various interactions following gauge-invariance Georgi et al. (1986); Jaeckel and Ringwald (2010); Ringwald (2014); Bauer et al. (2017); Brivio et al. (2017), in a manner similar to that of the QCD axion. However, unlike the QCD axion, their mass is generally unrelated to the decay constant, leading to more possible parameter space in axion experiments. These interactions are suppressed by the symmetry breaking scale, and can occur with gluons, photons, electroweak gauge bosons or fermions, and are being searched for in several experiments Choi et al. (2021); Giannotti (2023). In our study, we will consider such interactions with fermions carrying lepton number Dalla Valle Garcia et al. (2024); Domcke et al. (2020); Chun and Jung (2024), which, as we will see, has close connections with neutrino mass generation and leptogenesis Domcke et al. (2020); Chun and Jung (2024); Fong et al. (2023); Datta et al. (2024).
@@
Now, the common lore is that the axion field is initially static and misaligned at a non-zero field value, and later oscillates when its mass becomes comparable to the Hubble expansion rate of the Universe Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983). However, the axion field may instead be initially rotating in the complex field space Gouttenoire et al. (2021a); Li et al. (2014); Co et al. (2022a); Gouttenoire et al. (2021b); Harigaya et al. (2023a, b); Chung and Tadepalli (2024); Duval et al. (2024). Such an initial kick in the angular direction is usually realized at high-radial field values, through operators explicitly breaking the global symmetry. These dynamics at high radial field values have been crucial in cosmology, for e.g., in realizing Affleck-Dine baryogenesis Affleck and Dine (1985). Such setups have been explored further in the context of baryogenesis Co and Harigaya (2020), for modifying the dark matter abundance through the kinetic misalignment Co et al. (2020a, b), for cogenesis of both dark matter and the baryon asymmetry Co et al. (2021a, b) and also in several other related contexts Jeong and Takahashi (2013a, b); Higaki et al. (2014); Co et al. (2021c); Kawamura and Raby (2022); Co et al. (2021d, 2022b); Barnes et al. (2023); Co et al. (2023); Badziak and Harigaya (2023); Berbig (2024a); Chun and Jung (2024); Chao and Peng (2023); Barnes et al. (2024); Datta et al. (2024).
@@
In this paper, we explore the possibility of generating such a rotation of the axion, without delving into the above route, such that the radial mode still remains at the minima of its potential. This can take place if the height of the axion potential barrier keeps on decreasing, such that the axion can slide over,in a similar manner to Ricci reheating Opferkuch et al. (2019); Bettoni et al. (2022) (see also Refs. Figueroa and Byrnes (2017); Nakama and Yokoyama (2019)). Such an alternative was explored recently in Chun et al. (2024), with a time-dependent axion decay constant, realized through symmetry non-restoration. Here, we propose a scenario where an axion couples non-minimally to gravity, in a periodic form, thereby respecting the shift-symmetry of any pNGB action Ferreira et al. (2018). In contrast to Refs. Gouttenoire et al. (2021a); Li et al. (2014); Co et al. (2022a); Gouttenoire et al. (2021b); Harigaya et al. (2023a); Duval et al. (2024), our rotating axion always remains a spectator field. Because of its non-minimal coupling to gravity which depends on the cosmological background through the Ricci scalar $R$, the axion vacuum manifold evolves depending on the barotropic parameter $w$. While $w$ changes from 1-1 during inflation to +1+1 during kination (right after inflation), the tilt of the axion vacuum manifold flips in the opposite direction  [^1] Refs. Takahashi and Yin (2019); Huang et al. (2020) considered a similar flip without utilising the non-minimal coupling to gravity., generating an initial kinetic energy density for the axion. Interestingly, if the Universe happens to be kination-dominated after inflation, the potential barrier height being determined by the cosmological background, redshifts and decreases fast enough such that the axion can keep sliding across the barrier, and hence rotate in the complex field space.
@@
We utilise such a dynamics of the axion field to successfully generate both the baryon asymmetry and the dark matter abundance, dubbed as cogenesis. While the baryon asymmetry can be generated during the rotation through spontaneous baryogenesis Cohen and Kaplan (1987, 1988), the dark matter abundance can arise at a much later epoch, after the bare mass potential of the axion starts to dominate over its kinetic energy density. The observed value of baryon asymmetry constrains the non-minimal coupling $\xi$, while the dark matter abundance relates the axion decay constant to its mass. Importantly, because of the kination-dominated stiff era, the spectrum of the primordial gravitational waves (GWs) generated during inflation becomes blue-tilted for modes entering the horizon during kination. Such an increase in the GW amplitude tightly constrains our parameter space, because of an excess of GWs during Big Bang Nucleosynthesis (BBN).
@@
Throughout the paper we use natural units for which $\hbar=k_{B}=c=1$ and $8\pi G=m_{P}^{-2}$, with $m_{P}=2.43\times 10^{18}\,$GeV being the reduced Planck mass. The signature of the metric is positive.
@@ II The Framework
introduced in Ref. Ferreira et al. (2018); Salvio (2021); Ghoshal et al. (2023). In the above, $(\partial\phi)^{2}\equiv\partial_{\mu}\phi\,\partial^{\mu}\phi$ and
@@
When $|\phi|\ll f$, $1-\cos(\phi/f)]\simeq\frac{1}{2}(\phi/f)^{2}$. Then, the theory in Eq. (1) becomes
@@
which is reminiscent of the usual Lagrangian density of a scalar field with a non-minimal coupling to gravity [^2] Refs. Takahashi and Yamada (2015); Berbig (2024b) considered such a non-minimal coupling of the axion to gravity, where the latter anticipated a similar dynamics as in our scenario.. However, for a rotating axion the approximation $|\phi|\ll f$ is not valid ($0\leq\phi<2\pi f$), and we have to consider the full Lagrangian density in Eq. (1) without this approximation.
@@
In FRW spacetime $R=3(1-3w)H^{2}$, where $w$ and $H$ are the barotropic and the Hubble parameter, respectively. We consider that our axion is a spectator field in a non-oscillatory inflationary scenario, where inflation (with w=1w=-1) is followed by a period of kination (with $w=1$). We see, therefore, that during inflation, the axion effective potential is
@@
For the rolling of the axion to occur we need that $V_{\rm eff}^{\prime\prime}||>(\frac{3}{2}H)^{2}$, where $V_{\rm eff}^{\prime\prime}||=3\xi m_{P}^{2}H^{2}/f^{2}$. This condition suggests $\xi>\frac{3}{4}(f/m_{P})^{2}$, such that the range of $\xi$ is
@@
This condition also makes sure that during inflation the axion is driven towards the minimum expectation value $\phi=\pi\,f$. Because the axion is heavy during inflation, it does not undergo particle production. As a result, it does not introduce an isocurvature perturbation.
@@
[…]
@@
where $0<\lambda\lesssim 1$ is its self-coupling constant. To the above we add a temperature correction due to interaction with the thermal bath
@@
Assuming that the thermal bath is generated at the end of inflation (there are many mechanisms that do so naturally, e.g., by instant preheating Felder et al. (1999b); Dimopoulos et al. (2018) or Ricci reheating Dimopoulos and Markkanen (2018); Opferkuch et al. (2019); Bettoni et al. (2022)), we obtain an estimate of the maximum temperature $T_{\rm max}$ in the following way. After the end of inflation there is a period of kination, when the density of the Universe scales as $\rho\propto a^{-6}$, while radiation, which appears at the end of inflation, scales as $\rho_{r}\propto a^{-4}$. This means that
@@
This bound ensures that $|\Phi|\neq 0$ and the axion always exists. Considering GUT-scale inflation $H_{\rm end}\sim 10^{-5}m_{P}$. Then, because $T_{\rm reh}\gtrsim 10^{7}\,$GeV to avoid overproduction of gravitational radiation at BBN (cf. Eq. (49)), we obtain
@@
Before closing this section, we verify our claim of axion rotation above by numerically solving the equation of motion for the non-minimally coupled axion in a kination background. The equation of motion for $\theta=\phi/f$ is given by
$\ddot{\theta}+3H\dot{\theta}+V^{\prime}(\phi)/f=0\,$, (18)
where $V^{\prime}(\phi)/f=\frac{3\xi m_{P}^{2}H^{2}}{f^{2}}\sin\theta$. Now, after the end of inflation and flip of the axion vacuum manifold, $\theta$ is located at $\theta_{i}=-\pi$ as explained earlier. The axion receives an initial kick due to the tachyonic mass at the top of the potential. The magnitude of this quantum kick can be estimated as $\delta\phi=m/2\pi$ per Hubble time $\delta t=H^{-1}$ Felder et al. (2001), where $m^{2}=-V^{\prime\prime}(\theta_{i})$ indicates the tachyonic mass-squared at the top of the potential. Thus, the initial kinetic energy at $\theta_{i}=-\pi$ can be estimated as
$\frac{1}{2}f^{2}\dot{\theta}_{i}^{2}=\frac{1}{2}\left(\frac{\delta\phi}{\delta t}\right)^{2}\!\!\simeq\frac{3\xi}{8\pi^{2}}\left(\frac{H_{\rm end}}{f}\right)^{2}\!\!H_{\rm end}^{2}m_{P}^{2}\;,$ (19)
where we considered Eq. (7). In the limit of large $\xi$, i.e. $\xi\gg\left(f/m_{P}\right)^{2}$, the maximum velocity attained can be estimated as
$\frac{1}{2}f^{2}\dot{\theta}_{\rm m}^{2}$ $\simeq 6\xi m_{P}^{2}H_{\rm end}^{2}$ (20)
$\implies\dot{\theta}_{\rm m}$ $\simeq\frac{\sqrt{12\xi}m_{P}H_{\rm end}}{f}\leavevmode\nobreak\ ,$ (21)
where $H_{\rm end}$ denotes the Hubble scale at the end of inflation ($H_{\rm end}\simeq 10^{13}\,$GeV for GUT-scale inflation). [^3] To avoid excessive isocurvature perturbations, since the tachyonic kick is stochastic, we require $\dot{\theta}_{i}<\dot{\theta}_{\rm m}$. This translates into the bound $2f\gtrsim H/2\pi$, which is comparable to the bound discussed after Eq. (67). When $\xi$ is close to the lower limit $\left(f/m_{P}\right)^{2}$ (cf. Eq. (9)), the Hubble friction becomes comparable to the effective mass, and the axion is slowed down when it moves to the minimum. By this time, the potential energy also decreases, and the above estimate of the velocity $\dot{\theta}_{m}$ no longer holds.
[…]